2. Sketch the region of integration for SL to dz dy de and rewrite in the order dy dz dx.
Answer #1

Sol. s igo dz dy de 0 - 2. The order of integration is dz dy de i limits are a=0 to 54-y2 y=-2 to 2 a = o to i : 2= 54-yz CirRe- writing the limits. to 4-ze y = -√4-z2 z 0 to 2. a = o tol 1 2 4-92 . Integral is dy dz da. O 0-14-22